1986 explosion at the Chernobyl nuclear power plant in the former soviet union sent about 1000 kilograms of radioactive cesium-137 into the atmosphere. The function f(x)=1000(0.5)^x/30 describes the amount, f(x) in kilograms, of cesium-137 remaining in Chernobyl x yrs after 1986. If even 100 kilograms of cesium-137 remain in Chernobyl's atmosphere, the area is considered unsafe for humans. find f(80) and determine if Chernobyl will be safe to live in 2066.: 1986 explosion at the Chernobyl nuclear power plant in the former soviet union sent about 1000 kilograms of radioactive cesium-137 into the atmosphere. The function f(x)=1000(0.5)^x/30 describes the amount, f(x) in kilograms, of cesium-137 remaining in Chernobyl x yrs after 1986. If even 100 kilograms of cesium-137 remain in Chernobyl's atmosphere, the area is considered unsafe for humans. find f(80) and determine if Chernobyl will be safe to live in 2066.
Answer by josmiceli(2024)

f(x) = 1000*(.5)^(x/30)
f(80) = 1000*(.5)^(80/30)
f(80) = 1000*(.5)^(2.6667)
f(80) = 1000*.157486
f(80) = 157.486 answer
This is unsafe since it's more than 100 kg